A movie theater sells children's tickets for $7.00 and adult tickets for $9.00. The theater sold 120 tickets and earned $1,000.00. How many of each ticket did the theater sell?

Answer by mananth(16946)

Children's Tickets x
Adult tickets ------------ y
x + y = 120 .............1
7 x + 9 y =1000 .............2

multiply (1)by -9
Multiply (2) by 1
-9 x + -9 y = -1080
7x + 9 y = 1000
Add the two equations

-2 x = -80

/-2

x=40 Children's Tickets

plug value of x in (1)

x + y = 120
40+y = 120
y = 120-40
y = 80
y = 80 Adult tickets